Circular No. 5985

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

1993 RO, 1993 RP, 1993 SB, 1993 SC B. G. Marsden notes: “Ephemerides for 1993 RO, RP, SB, and SC,

derived from the ‘Pluto-like’ perihelion orbits on IAUC 5983, will be published on MPC 23544. They begin during June 17–July 17, when the objects should be near morning quadrature. Recovery of these mag 23–24 objects (one of which was observed over only a 2-day arc and two over a 6-day arc) is particularly important from the point of view of understanding the stability of the outer solar system, the nature of Pluto and the role of the Kuiper Belt, of which 1992 QB_{1} (for which there is also an ephemeris on MPC 23544) and 1993 FW (cf. IAUC 5939) are more obvious members, as perhaps also are 1994 ES_{2} (cf. MPEC 1994-H03) and 1994 EV_{3} (cf. MPEC 1994-H04). If the true orbits differ significantly from those on IAUC 5983, the ephemerides could be substantially in error—by perhaps 15′ if the objects are Neptune ‘Trojans’ (as is still a possibility, notably for 1993 RO) and by considerably more if the orbits are highly eccentric.”

COMET TAKAMIZAWA-LEVY (1994f)

Orbital elements from MPC 23482:

T = 1994 May 22.542 TT  ω = 61.^{o}635
Ω = 306.833 }2000.0
q = 1.35932 AU  i = 132.869
1994TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
May 8 20^{h}57.^{m}82 +29^{o}35.^{′}5 1.087 1.376 82.^{o}0 46.^{o}6 9.3
10 20 52.58 +32 48.1 1.050 1.372 83.5 47.0 9.2
12 20 46.37 +36 12.5 1.016 1.368 85.0 47.3 9.1
14 20 38.93 +39 47.7 0.986 1.365 86.3 47.6 9.0
16 20 29.96 +43 32.2 0.961 1.363 87.4 47.8 9.0
18 20 19.06 +47 23.5 0.940 1.361 88.4 48.0 8.9
20 20 05.67 +51 17.9 0.924 1.360 89.1 48.1 8.9
22 19 49.09 +55 10.4 0.914 1.359 89.6 48.1 8.8
24 19 28.41 +58 54.6 0.909 1.359 89.9 48.1 8.8
26 19 02.55 +62 22.4 0.910 1.360 89.9 48.1 8.8
28 18 30.41 +65 24.3 0.916 1.362 89.6 48.1 8.9
30 17 51.44 +67 49.4 0.928 1.364 89.1 48.0 8.9
June 1 17 06.40 +69 27.6 0.945 1.366 88.4 47.9 8.9

1994 May 6 Daniel W. E. Green